# Self-supervised Learning of Event-guided Video Frame Interpolation for Rolling Shutter Frames

### v1

@@ 5.2 Ablation Studies
Table 4: Ablation result on Gev-RS. $T$ indicates the count of time bins. $P$ indicates the use of pre-train optical flow model. $\mathcal{L}_{p}$ indicates the perceptual loss. $S_{h}$ and $S_{t}$ indicate the sample points for mask generation.
$\mathcal{L}_{gs2rs}$ $\mathcal{L}_{field}$ $\mathcal{L}_{rs2rs}$ $\mathcal{L}_{gs}$ PSNR SSIM
1 23.79 0.8076
2 23.91 0.8062
3 23.91 0.8091
4 25.08 0.8319
Table 5: Ablation of losses and supervised training for the 4 $\times$ interpolation on the Gev-RS dataset.
@@
Model complexity:Compared with other methods, our model stands out for its advantage in parameters and algorithmic complexity. Specifically, the parameter number of our framework is only one-fifth of EvUnRoll+TimeLens, as shown in Tab. 2. In addition, during the inference stage, our approach only needs to estimate DF once for generating all GS frames.In contrast, TimeLen and TimeReplayer require individual calculation of optical flow for each frame causing high computation cost.
6 Conclusion
In conclusion, this paper addressed the limitations of previous VFI methods by proposing a novel approach that considers the rolling shutter pattern commonly used in video capture.We introduced the displacement field that can effectively use the high temporal resolution property of events and model non-linear motion.Based on the displacement field, our framework realizes arbitrary frame rate VFI (e.g. 32×) as well as the reciprocal reconstruction between RS and GS frames in a self-supervised learning.We also employ two simulated datasets and collect a real-world dataset with spatially aligned events and RS RGB frames to support our cycle training.Experiments demonstrate the effectiveness of our method, both qualitatively and quantitatively.
Limitation and Future Work As the first attempt to employ a self-supervised approach for VFI based on RS images and events, our study has yielded promising results on simulated data. However, in the absence of quantitative metrics on real-world data, the efficacy of our method remains to be fully evaluated. In light of this limitation, future research should establish collaborations with camera manufacturers to capture simultaneous RS, GS, and events, thereby enabling robust and comprehensive evaluations for the development of innovative techniques for applications.

### v2

@@ V-B Efficiency Evaluation
TABLE IV: Ablation results on Gev-RS. $T$ indicates the count of time bins. $P$ indicates the use of pre-train optical flow model. $\mathbf{L}_{p}$ indicates the perceptual loss. $S_{h}$ and $S_{t}$ indicate the sample points for weight map.
$T$ P $\mathbf{L}_{p}$ GS-to-RS $\mathbf{L}_{p}$ RS-to-RS $S_{h}$ $S_{t}$ PSNR SSIM LPIPS
1 6 50 100 23.91 0.8091 0.0662
2 4 50 100 23.78 0.8019 0.0669
3 2 50 100 23.32 0.7928 0.0718
4 12 50 100 23.67 0.7960 0.0674
5 6 50 100 23.88 0.8082 0.0708
6 6 50 100 18.50 0.6225 0.1791
7 6 50 100 23.84 0.8134 0.0663
8 6 5 10 23.90 0.8097 0.0666
9 6 100 200 24.00 0.8109 0.0629
Figure 10: Comparison of the inference time of our method with TimeLens + EvUnRoll. We only count the time computing in the GPU, and I/O time is not included here. The x-axis is the magnification of the interpolated frame, from 3 times to 32 times. The y-axis represents time in milliseconds.
Figure 11: (a) The number of parameters required over time for event data and videos at different frame rates (fps). Parameters refer to the raw data size, such as for an $N$-frame video with a resolution of $H\times W$, the parameters are calculated as $N\times H\texttimes W$. For events, the parameters are the number of event signals multiplied by 4. The x-axis represents the recording time from 0 to 1 second, and the y-axis shows the amount of parameters required to store the data. The red and blue lines represent videos recorded at 128 fps and 24 fps, respectively, with a linear increase over time. The black line represents the average number of events, where the variability of the event recording depends on the scene. The different cyan-shaded areas show the range of this variation. (b) The distribution of event trigger times within 1 second. Specifically, in one second, the ratio of a pixel outputting 10 event signals is 0.16, the ratio of outputting 30 signals is about 0.05, and only less than 0.01 pixels output more than 70 signals.
Inference speed: To evaluate the efficiency of our method, we conducted the experiment on 32×32$\times$32 × interpolation with the resolution of $260\times 346$, as depicted in Fig. 10. Evidently, with the increment in interpolation factors, the inference time for EU+TL exhibits a substantial increase, whereas our method demonstrates only a marginal escalation from 312ms to 528ms. Notably, for a 32×32$\times$32 × interpolation, our technique requires merely 16ms to reconstruct each GS frame, which is approximately one-tenth of the time taken by EU+TL.
Transmission bandwidth analysis: We analyzed the correlation between the number of event activations and time using real data, and the results are visualized in the accompanying Fig. 11. In subfigure (a), the x-axis denotes time, while the y-axis signifies the average number of parameters needed per pixel. Notably, each activation requires one parameter for recording. The data suggests a linear increase in the number of activations over time. The median frequency of activations is approximately 17 times per second. Furthermore, 90% of the pixels activate fewer than 36 times, 80% activate less than 28 times, and 70% activate fewer than 24 times. Fig. 11 (b) illustrates the distribution of pixel activations within a 1-second duration. Notably, pixels activated 10 times constitute the majority at 16%, whereas those activated 50 times represent less than 10%.
By comparing the per-pixel params of event+fps 24 videos and fps 128 videos, we highlight the advantage of applying an event camera to reconstruct high-frame-rate video, especially with the consideration of limited transmission bandwidth. As illustrated in Fig. 11, the parameters of the event+fps 24 video are significantly smaller compared to those of the fps 128 video. This finding indicates that the event+fps 24 video is not only capable of providing sufficient spatial-temporal information to reconstruct high-frame-rate GS videos but also effectively mitigates the demand for transmission bandwidth.
@@
Model complexity: Compared with other methods, our model stands out for its advantage in parameters and algorithmic complexity. Specifically, the parameter number of our framework is only one-fifth of EvUnRoll+TimeLens, as shown in Tab. III. In addition, during the inference stage, our approach only needs to estimate DF once for generating all GS frames. In contrast, TimeLen and TimeReplayer require individual calculations of optical flow for each frame, causing high computation costs.
TABLE V: Ablation of losses and supervised training for the 4 $\times$ interpolation on the Gev-RS dataset.
$\mathbf{L}_{gs2rs}$ $\mathbf{L}_{field}$ $\mathbf{L}_{rs2rs}$ $\mathbf{L}_{gs}$ PSNR SSIM LPIPS
1 23.79 0.8076 0.0672
2 23.91 0.8062 0.0680
3 23.91 0.8091 0.0662
4 25.08 0.8319 0.0647
V-C Ablation Studies
We conduct the following ablation experiments to evaluate the effectiveness of our proposed modules in Gev-RS dataset [4]: E-RAFT Channel, Time bins and iters of E-RAFT are set to 15, 6, and 12 in the baseline case which uses pretrained E-RAFT model [55].
Loss function: Tab.V for each loss in Eq. 14. The 1st row represents the baseline, where solely the $\mathbf{L}_{gs2rs}$ loss (Eq. 10) is utilized, In the 2nd and 3rd rows, the displacement field loss (Eq. 3) and RS-to-RS loss (Eq. 13) are incorporated concurrently. As a result, increases are observed in the PSNR (+0.12) and SSIM (+0.0015). Also, when perceptual loss is added in our self-supervised setting (Tab. 4), we find it hard to obtain favorable outcomes.
Perceptual loss: Perceptual loss $\mathbf{L}_{p}$ is widely used in previous work[46, 3, 26, 39, 17]. We also perform experiments to study the effectiveness of the perceptual loss [72] by applying the perpetual loss on GS-to-RS reconstruction and RS-to-RS warping. Tab. IV (row 6-7) validates that introducing the perceptual loss can not boost the performance. Especially, applying perceptual loss only on the GS-to-RS reconstruction leads to the collapse of the displacement field and a worse performance. We argue that although the perceptual loss is applied in the compared method, it can not serve as the constraint in our self-supervised method.
Time bins: We conducted the experiments to evaluate the influences of different time bins on the displacement field estimation. As shown in Tab. IV (row 1-4), the settings with six-time bins obtain the best PSNR and SSIM scores. The explanation is that when the number of time bins goes up, the 3D deformation field will be divided into more segments bringing better performance because of more capacity of the non-linear motion. However, an excessive number of time bins can result in accumulated errors of motion and the degradation of performance. Fig. 6 shows warp weights corresponding to GS at different times. It can be observed that the nonlinear motion of the bus is estimated clearly.
Pretrained optical flow model: We evaluate the effect of the pre-train optical flow model by removing E-RAFT pre-train model. As shown in Tab. IV (row 1,5), the experiments with the pre-train model and without it demonstrate a similar performance. This indicates that our method can learn the optical flow without supervision.
Sample points of height and time of weight map: We investigate the impact of different sample points in height and time dimension during the estimation of the weight map, as shown in Fig. 4. Tab. IV (row 1,8-9) shows that the settings with different sample points have similar performance. This is because, for the setting with five height samples points and ten time samples points, each element has up to 50 samples points; thus, each element has ample samples to estimate the weight of warping.
VI Conclusion
In summary, we presented a novel method for reconstructing high-frame-rate (slow-motion) videos free from skew and jelly effects, thus enhancing the VR experience, which was previously constrained by consumer cameras hardware limitations. Our approach enables arbitrary frame rate video frame interpolation (e.g., 32×) and reciprocal reconstruction between RS and GS frames through self-supervised learning. Limitation and Future Work As the first attempt to employ a self-supervised approach for VFI based on RS images and events, our study has yielded promising results on simulated data. However, in the absence of quantitative metrics on real-world data, the efficacy of our method remains to be fully evaluated. Further research will consider how to combine event cameras with consumer cameras, e.g., UAVs, to collect aligned real-world RGB frames and events.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (NSF) under Grant No. 62206069 (affiliated with Guangzhou HKUST FYTRI), the MOE AcRF Tier 1 SSHR-TG Incubator Grant FY24 under Grant No. RSTG7/24, and the Open Project Program of State Key Laboratory of Virtual Reality Technology and Systems, Beihang University (No. VRLAB2025C04).